pooja195 (pooja195):

Solve using substitution. 5x + 12y = 146 x - 12y = -158

OpenStudy (abhisar):

First find the value of x in terms of y. \(\sf x+158=12y\\ \Rightarrow x=12y-158\)

OpenStudy (abhisar):

Now substitute this value of x in other equation.

pooja195 (pooja195):

5(12y-158)+ 12y = 146 60y-790+12y=146 72y-790=146

pooja195 (pooja195):

is that right so far?

OpenStudy (abhisar):

yes

pooja195 (pooja195):

72y-790=146 add 790 to both sides 72y=-644 this is the part i get stuck on

pooja195 (pooja195):

when i divide I get a fraction =/

OpenStudy (abhisar):

\(\sf 72y=146+790 \\ \huge y = \frac{146+790}{72}\)

pooja195 (pooja195):

oh whoops 936/72=13 yay!! i got :D

pooja195 (pooja195):

so y=13

pooja195 (pooja195):

wait now im confused where would i sub?

pooja195 (pooja195):

:(

OpenStudy (abhisar):

Yes, y =13. Now substitute this value of y in any of the original equations and calculate the value of x

pooja195 (pooja195):

x=12y−158 x=12(13)-158 x=156-158 x=-2?

TheSmartOne (thesmartone):

Correct :)

pooja195 (pooja195):

yay!